6) The rate constant for the first-order decomposition of N2O5 in the reaction 2N2O5(g) → 4NO2(g) + O2(g) is k=3.38 x 10-5 s-1 at 25°C. What is the half-life of N2O5? What will be the total pressure, initially 88.3 kPa for the pure N2O5 vapour, (a) 10 s, (b) 10 minutes after initiation of the reaction?
Answer #1

K= 3- 38x105 57 (t112) = 0.693 - 00693 half life 3- 38x155 1 a (t112= 20503 seconds answes. Po=88.3 kPa 6 t= io sec. p = po é